# Operationalizing Pluralistic Values in Large Language Model Alignment Reveals Trade-offs in Safety, Inclusivity, and Model Behavior

### v1

@@ Acknowledgements
We thank Michèle Wieland, Aysenur Kocak, Cheng Yu, Furkan Kadıoğlu, Jana Diesner, Shaghayegh Kolli and Nafiseh Nikeghbal for their valuable discussions and support throughout this project. Their insights contributed to shaping the study and improving the clarity of the work.
@@ A.2 Example of Gender-Focused Prompt-Response Pairs
Figure 6 display three example gender-related prompts collected from different datasets and Wizard-Vicuna-7B-Uncensored-GPTQ responses. The prompt-response pairs were translated into German using DeepL (2024). We selected gender-related prompts to ensure all participants evaluated responses within the same thematic domain, controlling for topic variation and isolating differences in individual judgment rather than content effects.
@@ A.3 Definitions of Alignment Dimensions
The following definitions were shown at the start of the survey to ensure consistent interpretation of each alignment dimension. Participants received the definitions directly in the survey interface (without additional examples) and were also provided with a link to our full paper containing extended explanations. To avoid language mixing, we deployed two parallel survey platforms: an English version and a German version. Participants were routed automatically based on their Prolific demographic information, and each participant saw the entire survey-definitions and instructions—in a single language only.
@@ A.4 Detailed Recruitment and Data Quality Procedures
We recruit participants from Prolific, a crowdsourcing platform designed for academic research  (Prolific 2025). Each participant rated 5 randomly drawn prompt-response pairs from a pool of 1,762. The total cost for participant compensation across both countries was approximately \$6,000 USD. We implement multiple quality control measures to ensure reliable data:
@@
  • Response pattern analysis: After data collection, we conduct post-hoc validation to identify participants who may have rushed their responses. This includes detecting uniform response patterns where participants select the same rating (e.g., “Strongly Agree”) across all five dimensions for multiple prompt-response pairs, indicating insufficient engagement with the content rather than genuine evaluation.
@@ A.5 Limitations and Future Work
Our dataset includes participants from the US and Germany, limiting cultural scope; we also recognize that both countries represent predominantly WEIRD (Western, Educated, Industrialized, Rich, and Democratic) cultures, and suggest that future work be conducted across a more diverse set of countries and cultures. We standardized political ideology using a five-point scale and harmonized ethnicity categories across countries, while acknowledging inherent cross-national differences. To ensure sufficient sample sizes, we collapsed adjacent ideological groups (e.g., “conservative” and “rather conservative”). (See Supplementary Table 3 for participant demographics) Still, conservatives, gender minorities, and older adults remained underrepresented, reflecting common online recruitment challenges. Future work should expand geographic and demographic diversity.
@@ B.2 Complete Model and Experimental Specifications
All models used in this study were available on Hugging Face throughout the fine-tuning period (March–September 2025), ensuring stable and reproducible training. The selection captures variation in size, objectives, and safety properties to support comparative analysis across architectures. Table 9 lists the models included.
@@ B.3 Evaluation Methodology
We evaluate alignment performance using OpenAI’s GPT-4o-mini API to score model responses on Toxicity (continuous 0.0-1.0 scale) and Emotional Awareness dimensions (binary 0-1). We use two different scales because toxicity, following the continuous scoring framework of (Jigsaw and Google ), varies by degree, whereas emotional awareness, consistent with Sharma et al. (2020), is expressed in threshold-like categories and is therefore more reliably captured as a binary judgement. To validate this automated approach, we conduct a reliability study where two independent human experts assess 50 randomly selected responses using identical criteria. We achieve an average agreement rate of 0.85 between human evaluators and automated scores across both dimensions, confirming the reliability of our evaluation approach. Our automated evaluation employs GPT-4o-mini with temperature=0 for consistent scoring. Figure 9 shows the prompts used for each alignment dimension.
@@ B.4 Alignment Optimization Methods
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}\left[\log\sigma\left(s_{\theta}(y^{+};x)-s_{\theta}(y^{-};x)\right)\right],$ (2)
@@
To optimize two objectives simultaneously (e.g., toxicity and emotional awareness), we pool all accept/reject pairs-”Accept” refers to the response rated higher by participants on the target dimension, and ”reject” is the lower-rated response- from both dimensions into one dataset and apply the same loss:
$\mathcal{L}_{\text{multi-DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}\big[\log\sigma\big(s_{\theta}(y^{+}\mid x)\\ -s_{\theta}(y^{-}\mid x)\big)\big]$ (10)
@@
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}\left[\log\sigma\left(s_{\theta}(y^{+})-s_{\theta}(y^{-})\right)\right].$ (11)

### v2

@@ Acknowledgements
We thank Michèle Wieland, Aysenur Kocak, Cheng Yu, Furkan Kadıoğlu, Jana Diesner, Shaghayegh Kolli and Nafiseh Nikeghbal for their valuable discussions and support throughout this project.
@@ A.2 Example of Gender-Focused Prompt-Response Pairs
Figure 6 displays three example gender-related prompts collected from different datasets and Wizard-Vicuna-7B-Uncensored-GPTQ responses. The prompt-response pairs were translated into German using DeepL (2024). We selected gender-related prompts to ensure all participants evaluated responses within the same thematic domain, controlling for topic variation and isolating differences in individual judgment rather than content effects.
@@ A.3 Definitions of Alignment Dimensions
The following definitions were shown at the start of the survey to ensure consistent interpretation of each alignment dimension. Participants received the definitions directly in the survey interface (without additional examples) and were also provided with links to the source papers that define each dimension. To avoid language mixing, we deployed two parallel survey platforms: an English version and a German version. Participants were routed automatically based on their Prolific demographic information, and each participant saw the entire survey-definitions and instructions—in a single language only.
@@ A.4 Detailed Recruitment and Data Quality Procedures
We recruit participants from Prolific, a crowdsourcing platform designed for academic research  (Prolific 2025). Data collection took place between September 2024 and January 2025. Each participant rated 5 randomly drawn prompt-response pairs from a pool of 1,762. The total cost for participant compensation across both countries was approximately \$6,000 USD. We implement multiple quality control measures to ensure reliable data:
@@
  • Response pattern analysis: After data collection, we conduct post-hoc validation to identify participants who may have rushed their responses. This includes detecting uniform response patterns where participants select the same rating (e.g., “Strongly Agree”) across all five dimensions, indicating insufficient engagement with the content rather than genuine evaluation.
@@ A.5 Limitations and Future Work
Our dataset includes participants from the US and Germany, limiting cultural scope; we also recognize that both countries represent predominantly WEIRD (Western, Educated, Industrialized, Rich, and Democratic) cultures, and suggest that future work be conducted across a more diverse set of countries and cultures. We standardized political ideology using a five-point scale and harmonized ethnicity categories across countries, while acknowledging inherent cross-national differences. To ensure sufficient sample sizes, we collapsed adjacent ideological groups (e.g., “conservative” and “rather conservative”) – see Supplementary Table 3 for participant demographics. Still, conservatives, gender minorities, and older adults remained underrepresented, reflecting common online recruitment challenges. Future work should expand geographic and demographic diversity.
@@ B.2 Complete Model and Experimental Specifications
All models used in this study were available on Hugging Face throughout the fine-tuning period (March–September 2025), ensuring stable and reproducible training. The selection captures variation in size, objectives, and safety properties to support comparative analysis across architectures. Table 9 lists the models included. Model-level results, including full breakdowns for every architecture, are provided in our public repository https://github.com/DALIAALISIDDIG/AlignCure.
@@ B.3 Evaluation Methodology
We evaluate alignment performance using OpenAI’s GPT-4o-mini API to score model responses on Toxicity (continuous 0.0-1.0 scale) and Emotional Awareness dimensions (binary 0-1). We use two different scales because toxicity, following the continuous scoring framework of (Jigsaw and Google 2025), varies by degree, whereas emotional awareness, consistent with Sharma et al. (2020), is expressed in threshold-like categories and is therefore more reliably captured as a binary judgement. To validate this automated approach, we conduct a reliability study where two independent human experts assess 50 randomly selected responses using identical criteria. We achieve an average agreement rate of 0.85 between human evaluators and automated scores across both dimensions, confirming the reliability of our evaluation approach. Our automated evaluation employs GPT-4o-mini with temperature=0 for consistent scoring. Figure 9 shows the prompts used for each alignment dimension.
@@ B.4 Alignment Optimization Methods
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}[\log\sigma(s_{\theta}(y^{+};x)-s_{\theta}(y^{-};x))].$ (2)
@@
To optimize two objectives simultaneously (e.g., toxicity and emotional awareness), we pool all accept/reject pairs—“accept” refers to the response rated higher by participants on the target dimension, and “reject” is the lower-rated response—from both dimensions into one dataset and apply the same loss:
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}[\log\sigma(s_{\theta}(y^{+};x)-s_{\theta}(y^{-};x))].$ (10)
@@
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}[\log\sigma(s_{\theta}(y^{+};x)-s_{\theta}(y^{-};x))].$ (11)